# Revisiting Differentially Private Hyper-parameter Tuning

### v2

@@ 6.1 Stronger Audit via Reduction
TABLE IV: Comparison of various privacy results for $\mathcal{H}$ (Algorithm 2). All results are in $(\varepsilon,\delta=10^{-5})$-DP form. Value for $\varepsilon_{L}$ (the lower bound) is the average of 10 runs with different seeds for simulating the game shown in Figure 7. $\varepsilon^{O}_{\mathcal{H}}$ is our improved upper bound (Algorithm 4 with the base algorithm is some $\mu$-GDP) and $\varepsilon_{\mathcal{H}}^{P}$ is the privacy upper bound derived by previous work [42]. $\varepsilon_{L}$, $\varepsilon^{O}_{\mathcal{H}}$ and $\varepsilon_{\mathcal{H}}^{P}$ in each cell are all computed based on the same base algorithm’s privacy budget $\varepsilon_{B}$ shown in leftmost column. Note that $\tau$ is the sampling ratio with $N=10^{3}$ total iteration, and $\eta,\nu$ is the parameter for TNB distribution (shown in Appendix A.2; $\mathbb{E}_{\xi}$ is its expectation) input to $\mathcal{H}$. We can see that our result is better under all setups.
2) Instantiate the score function. Based on our above reduction, to serve the distinguishing purpose, a model obtained by DP-SGD can be “treated” as a real number sampled from univariate Gaussian or its shifted counterpart corresponding to $X$ or $X^{\prime}$ was used. With our reduction, the order induced by the score function $g$ is now over $\mathbb{R}$. To have stronger audit results in our idealized attack, we need to instantiate the worst-case score function, and our Theorem 2 tells us one-to-one mapping score function $g$ is the worst case necessarily.
However, Theorem 2 remains silent on the specific analytical form of $g$ in the worst case. We will discuss the hardness of finding the specific worst-case $g$ in Section 6.2. Based on our above reduction, there can be infinitely many one-to-one mapping functions $\mathbb{R}\rightarrow\mathbb{R}$; for implementation purposes, we now fix a score function $g(x)=x$, i.e., we take $g$ is strictly increasing and all strictly increasing functions induces the same order over $\mathbb{R}$ regardless of its analytical form. To summarise, we present the distinguishing game of our reduced case in Figure 7. Our further audit experiment in this section is to simulate such a game many times and conclude the lower bound in accordance with Section 2.2.
6.2 Lower Bound & Privacy Results Comparison
We present the lower bound $\varepsilon_{L}$, our improved privacy results $\varepsilon^{O}_{\mathcal{H}}$ and previous privacy results $\varepsilon_{\mathcal{H}}^{P}$ [42] in Table IV. Fixing $\delta_{\mathcal{H}}=10^{-5}$, we can see that our privacy result $\varepsilon_{\mathcal{H}}^{O}$ is better than previous result $\varepsilon_{\mathcal{H}}^{P}$ under all setups. We can also see that $\varepsilon_{L}$ is higher than $\varepsilon_{B}$, confirming the very action of selecting the best does leak additional privacy beyond the base algorithm’s privacy budget, similar to our finding in Section 4.5 (ETCV). However, we must emphasize that such a phenomenon is only observed under a strong adversary setup.
Lessons Learned and Open Problems. As shown in Section 5.3, our improved result is indeed tight for general $\xi$ in terms of how much $(\varepsilon_{\mathcal{H}},\delta_{\mathcal{H}})$-DP $\mathcal{H}$ satisfies. However, we still see a noticeable gap between our result $\varepsilon^{O}_{\mathcal{H}}$ and the lower bound $\varepsilon_{L}$ derived by our idealized attack. why does this happen? Answering this question will lead us to a deeper understanding of how selection leaks privacy.
@@ 6.2 Lower Bound & Privacy Results Comparison
$g(x)=\left\{\begin{array}[]{ll}x,\hskip 15.93347pt\text{for $x\in[-\infty,0)\bigcup(1,\infty]$}\\ 1-x,\text{for $x\in[0,1]$}\\ \end{array}\right.$ (27)
@@
Choosing some $g$ arbitrarily and performing the attack will likely end up with sub-optimal attacks (smaller/weaker lower bounds). This is possibly the reason why we still see a gap between $\varepsilon^{O}_{\mathcal{H}}$ and $\varepsilon_{L}$ even in our idealized attack. This poses a critical question for the adversary, i.e., which $g$ should the adversary choose to elicit more privacy leakage? Another equally important question is: does the worst-case $g$ depend on specific $\xi$? Answering the above questions is non-trivial and requires future investigations.
7 Conclusion
We aim to gain a deeper understanding of current theoretical v.s. practical privacy for the private selection problems. Initially, we give examples showing that the current generic bound for private selection is indeed tight in general. However, when we consider a white-box setting, i.e., the hyper-parameter tuning problem, such tightness no longer applies. Substantiating this assertion, we first conduct empirical studies for private hyper-parameter tuning under various privacy audit setups. Our findings reveal that, in certain practical scenarios, a weak adversary can only acquire restricted sensitive information. The derived empirical privacy bound for the strongest adversary still sees a noticeable gap from the generic upper bound.
We then take further steps to analyze privacy in hyper-parameter tuning and derive better privacy results by modeling the base algorithm’s privacy with finer resolution ($f$-DP). The improvement is due to the distinct properties of the base algorithm (DP-SGD). Our result is better under various settings than the current generic bound. In addition, our analysis generalizes, contrasting with previous work, where their results are only easily applicable to specific parameter configurations.
@@ Appendix A Content for reference
A.1 Neyman–Pearson Lemma
@@
A.2 Privacy Results by Papernot et al.
@@
Let $k$ in Algorithm 1 follows TNB distribution $\xi_{\eta,\nu}$. If the base algorithm satisfies $(\alpha,\gamma)$-RDP and $(\alpha^{\prime},\gamma^{\prime})$-RDP, Algorithm 1 satisfies $(\alpha,\hat{\gamma})$-RDP where
@@
A.3 Used Datasets and Experimental Details
Our implementation is provided at an anonymous link111 https://anonymous.4open.science/. We use four image datasets in our experiments. FASHION [55], MNIST [28], CIFAR10 [26] and SVHN [40]. All of our experiments are conducted under privacy parameter $\delta=10^{-5}$. The number of repeating/simulation times in an audit experiment is $2,000$. The error bar results from taking the min., max., and avg. for three trials. To efficiently audit the hyper-parameter tuning and reduce the simulation burden, we only fetch 5,000 data examples from the original training datasets, and we set the sampling rate to be 1, i.e., full gradient descent. We set the TNB distribution [42] with parameter $(\eta=0,\nu=10^{-2})$. We use the ResNet [22] as the neural network in our experiments. Details for network setup can be found in our implementations at our anonymous link. Our audit experiment consumes $>4000$ GPU hours and is conducted over 20 GPUs in parallel.
Hyperparameter candidates setup. To run the audit experiments, we need to set the candidates inside $\Omega$ in Algorithm 2. We hold the clipping threshold $C$, learning rate $\mathbf{lr}$, and the number of total iterations $N$ as the hyperparameters to be tuned. To form each candidate inside $\Omega$, we randomly sample a value to determine $C$, $\mathbf{lr}$ and $N$. Details can be found at our implementation. All candidates have the same privacy budget according to our problem formulation.
@@ A.3 Used Datasets and Experimental Details
Figure 8: The output’s score histogram under all setups.
A.4 More Experimental Results
Measuring the score distribution. To confirm the robustness mentioned above. A straightforward way is to inspect the scores of the output of $\mathcal{H}$ (Algorithm 2). We measure various training datasets, and results are presented in Figure 8. We process the scores by shifting so all scores are $<0$. Our result clearly shows no significant difference between the score distributions when $\mathbf{z}$ is included ($X^{\prime}$ is used) and when it is not included ($X$ is used) in training. In Figure 8, we again obverse the “indistinguishability” between cases where $\mathbf{z}$ is included and cases where $\mathbf{z}$ is not.
@@ B.2 Proof of Theorem 2
Similarly, the distribution of $F^{\prime}_{k,g^{*}}$ corresponding to $P^{\prime}$ has the same form (just replace $\Lambda_{S}^{k}$ by $\bar{\Lambda}_{S}^{k}$). We now compute the RDP quantity $\mathcal{D}_{\alpha}(F_{k,\hat{g}}||F^{\prime}_{k,\hat{g}})$ and $\mathcal{D}_{\alpha}(F_{k,g^{*}}||F^{\prime}_{k,g^{*}})$. We aim to show that the RDP value under non-one-to-one mapping $\hat{g}$ is smaller than that under its one-to-one mapping counterpart. We group the sub-terms of RDP calculation. Let
$\hat{T}_{\{a,c,e\}}=\sum_{i\in\{a,c,e\}}\left(\frac{F_{k,\hat{g}}(i)}{F^{\prime}_{k,\hat{g}}(i)}\right)^{\alpha}F^{\prime}_{k,\hat{g}}(i)$
$\hat{T}_{\{b,d\}}=\sum_{i\in\{b,d\}}\left(\frac{F_{k,\hat{g}}(i)}{F^{\prime}_{k,\hat{g}}(i)}\right)^{\alpha}F^{\prime}_{k,\hat{g}}(i)$
$\hat{T}_{\{f\}}=\left(\frac{F_{k,\hat{g}}(f)}{F^{\prime}_{k,\hat{g}}(f)}\right)^{\alpha}F^{\prime}_{k,\hat{g}}(f)$
@@ B.3 Proof of Theorem 3
[…]

### v3

@@ VI-B Privacy Results Comparison
TABLE IV: Comparison of privacy bounds for $\mathcal{H}$ (Algorithm 2). All values are in $(\varepsilon,\delta=10^{-5})$-DP form. $\varepsilon_{L}$ is the empirical lower bound obtained by simulating the distinguishing game in Figure 6. $\varepsilon_{\mathcal{H}}^{O}$ is our improved analytical upper bound. $\varepsilon_{\mathcal{H}}^{P}$ is the upper bound from prior work [42].Each row corresponds to a different sampling ratio $\tau$, with total iterations fixed at $N=10^{3}$. The parameters $\eta$ and $\nu$ define the TNB distribution used to generate the number of runs in $\mathcal{H}$ (details in Appendix A-B), and $\mathbb{E}_{\xi}$ is the expected number of runs under this distribution.
$\varepsilon_{B}$ $\varepsilon^{O}_{\mathcal{H}}$ Previous $\rightarrow$ Ours
MNIST FMNIST CIFAR10 SVHN
11 $1.83$ $0.921$ $\rightarrow$ $0.934$ $0.768$ $\rightarrow$ $0.793$ $0.412$ $\rightarrow$ $0.448$ $0.636$ $\rightarrow$ $0.661$
22 $3.43$ $0.942$ $\rightarrow$ $0.956$ $0.779$ $\rightarrow$ $0.802$ $0.467$ $\rightarrow$ $0.486$ $0.706$ $\rightarrow$ $0.745$
44 $6.69$ $0.951$ $\rightarrow$ $0.958$ $0.791$ $\rightarrow$ $0.817$ $0.504$ $\rightarrow$ $0.531$ $0.762$ $\rightarrow$ $0.786$
TABLE V: Testing accuracy comparison under differentially private hyper-parameter tuning. The TNB setup for our method is $(\eta=0,\nu=10^{-3})$ ($\mathbb{E}_{\xi}\approx 144$). To achieve roughly the same privacy result using previous method [42], the setup should be $(\eta=0,\nu=10^{-2})$ ($\mathbb{E}_{\xi}\approx 21$). $\varepsilon^{O}_{\mathcal{H}}$ is our improved result for private hyper-parameter tuning. For all experiments, $\delta=10^{-5}$.
VI-C Lessons Learned and Open Problems
As shown in Section V-C, our improved result is indeed tight for general $\xi$ in terms of how much $(\varepsilon_{\mathcal{H}},\delta_{\mathcal{H}})$-DP $\mathcal{H}$ satisfies. However, we still see a noticeable gap between our result $\varepsilon^{O}_{\mathcal{H}}$ and the lower bound $\varepsilon_{L}$ derived by our idealized attack. why does this happen?
@@ VI-C Lessons Learned and Open Problems
[…]
@@
Choosing some $g$ arbitrarily and performing the attack will likely end up with sub-optimal attacks (smaller/weaker lower bounds). This is probably the reason why we still see a gap between $\varepsilon^{O}_{\mathcal{H}}$ and $\varepsilon_{L}$ even in our idealized attack. Reasoning on such issues is non-trivial, and it poses the following questions worthy of investigation:
1) which $g$ should the adversary choose to elicit more privacy leakage? 2) does the worst-case $g$ depend on specific $\xi$? 3) How to quantify the exact trade-off between privacy leakage and $\xi$ which governs the utility? Answering the above questions requires non-trivial efforts, which we hold as meaningful future directions.
VII Conclusion
We study how selection leaks privacy. Initially, we give examples showing that the current generic bound for private selection is indeed tight in general. Still, it is not tight for a white-box setting, i.e., the hyper-parameter tuning problem. Substantiating this assertion, we first audit the privacy of hyper-parameter tuning under various settings; the derived empirical privacy lower bound under the strongest adversary still sees a noticeable gap from the generic upper bound.
We then provide an in-depth study of deriving better privacy upper bound by modeling the base algorithm’s privacy with finer resolution ($f$-DP). The improvement is due to the distinct properties of the base algorithm (DP-SGD). Our result allows trying many more hyper-parameter candidates while consuming less private budget. Our analysis also generalizes, contrasting with previous work, which remains unknown how to adapt to general parameter setups.
VIII Ethics Considerations
This paper is on refining the privacy bound for differentially private protocols, not on privacy attacks. The privacy audit experiments conducted are to conclude a privacy lower bound, not to launch some real-world privacy attacks. All analyses and experiments are conducted using publicly available datasets to minimize privacy risks. Our study aims to strengthen differential privacy protections in hyper-parameter tuning by improving the analysis rather than exploiting any weaknesses.
@@ Appendix A Content for reference
A-A Neyman–Pearson Lemma
@@
A-B Privacy Results by Papernot et al.
@@
[…]
@@
A-C Used Datasets and Experimental Details
Our implementation is provided at an anonymous link [^2] https://github.com/zihangxiang/PrivateHyperparameterTuning.git. We use four image datasets in our experiments. FASHION [55], MNIST [28], CIFAR10 [26] and SVHN [40]. All of our experiments are conducted under privacy parameter $\delta=10^{-5}$. The number of repeating/simulation times in an audit experiment is $2,000$. The error bar results from taking the min., max., and avg. for three trials. To efficiently audit the hyper-parameter tuning and reduce the simulation burden, we only fetch 5,000 data examples from the original training datasets, and we set the sampling rate to be 1, i.e., full gradient descent. We set the TNB distribution [42] with parameter $(\eta=0,\nu=10^{-2})$. We use the ResNet [22] as the neural network in our experiments. We use Adam as the default optimizer. The computational burden is significant: our audit experiment consumes >4000>4000 GPU hours and is conducted over 20 GPUs in parallel.
Hyperparameter candidates setup. To run the audit experiments, we need to set the candidates inside $\Omega$ in Algorithm 2. We hold the clipping threshold $C$, learning rate $\mathbf{lr}$, and the number of total iterations $N$ as the hyperparameters to be tuned. To form each candidate inside $\Omega$, we randomly sample a value to determine $C$, $\mathbf{lr}$ and $N$. All candidates have the same privacy budget according to our problem formulation.
@@ B-B Proof of Theorem 2
[…]
@@ B-C Proof of Theorem 3
Define the privacy loss random variable for a pair of adjacent dataset $X,X^{\prime}$ to a private mechanism $\mathcal{M}$ as $L_{1}=\log\frac{\mathcal{M}(X)(o)}{\mathcal{M}(X^{\prime})(o)}$ where $o\sim\mathcal{M}(X)$ $\mathcal{M}$ is $(\varepsilon,\delta)$-DP or $f_{\varepsilon,\delta}$-DP if and only if